Four moles of carbon monoxide (CO) has 48 grams of carbon (C) and 64 grams of oxygen (O). What is the percent composition of carbon in carbon monoxide?
1
Expert's answer
2018-06-06T07:36:11-0400
Answer:
m (CO) = n*M, where n is the amount of substance, mole, M is the molar mass, g/mole
m (CO) = 4*28 = 112 g
% C = m (C)*100% / m (CO) = 48*100 % / 112 = 42.8 %